# Grounded Test-Time Adaptation for LLM Agents

### v3

@@ 4.1 Setup
We evaluate our approach on three realistic agentic benchmarks: BFCLv3 (patil_berkeley_nodate), WebArena (zhou_webarena_2024), and Tau-Bench (yao_tau-bench_2024), which together cover both web navigation and conversational function calling tasks. As for baselines, we evaluate open-source Qwen2.5-14B-Instruct (qwen_qwen25_2025) model and closed-source GPT4.1. On WebArena, we additionally compare to the World-model-augmented (WMA) agent from (chae_web_2025) as a baseline. This approach first collects exploration demonstrations, then trains a Llama-3.1-8B-Instruct (grattafiori_llama_2024) model to predict the next state given an action (i.e., a learned world model). During evaluation, this world model is used to provide next-state predictions to the agent on-the-fly, helping the agent reason about the consequences of its actions. In this comparison, we use the same GPT4o-mini as the original WMA experiments. Aside from the parametric and non-parametric experiments, we provide ablation studies in section 4.3.
For all non-parametric test-time adaptation experiments, we leverage GPT4.1 as an exploration policy to extract environmental dynamics. For each environment, we conduct $10$ exploration episodes to gather a diverse set of dynamics. For environment dynamics filtering, we apply o3222https://platform.openai.com/docs/models/o3 to remove repetitive and trivial environment dynamics. We also include ablation studies on the number of exploration episodes and the effect of filtering in section 4.3.
@@
The WebArena benchmark consists of five self-hosted websites and a total of 812 tasks. Many tasks require interacting across multiple websites, which poses additional challenges for LLM-based agents. For evaluation, WebArena employs a string matching based mechanism to assess correctness, providing more stable and objective results compared to benchmarks using LLMs to score, such as (he_webvoyager_2024; gou_mind2web_2025). For non-parametric test-time adaptation, we first synthesize $10$ personas using prompt instruction LABEL:lst:web_persona_synthesis for each website using GPT4o (openai_gpt-4_2024) and with seed personas from NNetNav (murty_nnetnav_2025). Then for each persona, we roll-out a GPT4.1 agent to explore the environment guided by the persona description, with the exploration prompt instruction LABEL:lst:web_exploration derived from NNetNav, and with a maximum exploration step of 30. To extract environment dynamics from interaction logs, we summarize each state transition pair (observation, action taken, new observation) with prompt instruction LABEL:lst:web_dynamics_extraction, and concatenate all summarized state transition pairs to obtain the list of environmental dynamics $E=\{e\}_{i=1}^{M}$ for each website. Finally, we utilize the o3 model to remove repeative and trivial environment dynamic entries from the list via prompt instruction LABEL:lst:web_filter_dynamics and this results in $E_{\text{clean}}$. We provide all prompts we use in appendix A. With parametric test-time adaptation, we reset the adaptation vector after each episode and we use a learning of 0.1 and training step of 1 for all experiments. During evaluation, we use the text-based accessibility tree as the observation type and run experiments using BrowserGym [^3] https://github.com/ServiceNow/BrowserGym and AgentLab [^4] https://github.com/ServiceNow/AgentLab with a maximum step of $30$.
Table 2: Main results on WebArena, WebVoyager, BFCLv3, and Tau-Bench benchmarks. We report task success rates (%) for each model and adaptation method. For Tau-Bench, we average runs across 5 seeds and use a custom more stable codebase. Both parametric (PA) and non-parametric (NPA) test-time adaptation improve performance. ”N/A” indicates not applicable—AWM requires training a web-specific world model on state transitions, which is not applicable to Tau-Bench (conversational only) or BFCLv3 (no explicit web-like state transition data); Non-parametric adaptation does not operate in conversational Tau-Bench as there is no explicit and fixed state transition rules in the environment.
@@
GPT 4.1 30.0 55.5 - -
GPT 4.1 (+ NP) 35.0 (+5.0) 64.0 (+8.5) N/A N/A
GPT 4o-mini 12.0 - - -
GPT 4o-mini (+ WMA) 13.5 (+1.5) N/A N/A N/A
GPT 4o-mini (+ NP) 18.0 (+6.0) - N/A N/A
@@
Qwen2.5-14B-Instruct (+ PA) 18.0 (+1.0) 20.0 (+1.5) 25.2 (+3.6) 44.9 (+1.6)
Qwen2.5-14B-Instruct (+ NPA) 20.0 (+3.0) 22.0 (+3.5) N/A N/A
@@
BFCLv3 (patil_berkeley_nodate) is a benchmark with various realistic APIs to evaluate the functions calling abilities of LLM. It contains $8$ different API domains. We provide the number of APIs for each environment in table 8. We choose the multi-turn-base split to assess the agent’s multi-turn function calling capabilities. This split is multi-turn and multi-step, meaning that instead of just handling a single tool call in a one-off request, models have to carry out sequences of function calls across multiple dialogue turns. BFCLv3 uses state-based evaluation (checking the system’s resulting state) and response-based evaluation (checking whether the call sequences are valid and minimally sufficient) to assess correctness, therefore being stable and deterministic. As for non-parametric test-time adaptation, we synthesize $10$ exploration tasks with GPT4o conditioned on the function documentations for each domains via prompt instruction LABEL:lst:bfcl_exploration_synthesis. Concretely, we instruct LLM to think about what functions it wants to explore in order to find out unpredictable outcomes or interesting behaviors, given the function documentations. We use a GPT4.1 agent to execute the exploration task and we collect interaction logs for environment dynamics extraction. During exploration, we ask the exploration agent to decide if steps taken so far are sufficient to answer the exploration task with prompt instruction LABEL:lst:bfcl_exploration_user_prompt. The environment dynamics are extracted using prompt instruction LABEL:lst:bfcl_extract_dynamics. We follow the same procedure as in WebArena to remove repetitive and trivial environment dynamics with prompt instruction LABEL:lst:bfcl_filter_env_dynamics. For parametric test-time adaptation, we reset adaptation vector at each turn (when a new user query arrives).
@@
Tau-Bench (yao_tau-bench_2024) is a realistic function calling benchmark that simulates real-world service agents interacting with an LLM-simulated user. It comprises two domains: airline (50 tasks) and retail (115 tasks). The airline domain is generally more challenging than retail, as it features a much larger database, increasing the complexity of the tasks. Tau-bench’s use of an LLM-simulated user agent introduces significant variance in evaluation results, as the simulated human may drift from the original instruction or be influenced by the agent’s responses over multiple conversational turns. To address this, we adopt the improved evaluation codebase from (prabhakar_apigen-mt_2025), which substantially reduces reward variance and improves evaluation fidelity. Specifically, we (1) fix incorrect labels, (2) set the temperature of the LLM-simulated user agent to 0 for deterministic responses, and (3) employ a Best-of-N (N=5) sampling strategy combined with a self-critique mechanism, allowing the simulated user to better adhere to the task instruction and avoid being misled by the agent’s outputs. This protocol has been shown to improve consistency in agent performance evaluation across multiple trials according to (prabhakar_apigen-mt_2025). Tau-Bench consists solely of API-based tasks without observable state transitions, meaning there are no environment dynamics to extract. As a result, non-parametric adaptation, which relies on extracting and leveraging such dynamics is not applicable to this benchmark. Therefore, we only evaluate parametric test-time adaptation method on Tau-Bench. To further reduce variance, we report results averaged over 5 random seeds in table 2.
@@ 4.2 Results and Analysis
Table 3: Success rates (%) on the WebArena benchmark across different websites and models. Both parametric adaptation (PA) and non-parametric adaptation (NPA) improve performance over the base models (GPT4.1, GPT4o-mini and Qwen2.5-14B-Instruct) and with non-parametric adpatation (NPA) surpassing WMA on GPT4o-mini. NPA improves success rate substantially on complex multi-site split.
@@
Model Gitlab Shopping Admin Shopping Reddit Map Multi-site Average
GPT4.1 48.0% 31.0% 37.0% 29.0% 28.0% 2.0% 30.0%
GPT4.1 (+NPA) 45.0% 27.0% 33.0% 42.0% 31.0% 23.0% 35.0%
GPT4o-mini 9.0% 12.0% 18.0% 8.0% 12.0% 12.0% 12.0%
GPT4o-mini (+WMA) - - - - - - 13.5%
GPT4o-mini (+NPA) 18.0% 14.0% 26.0% 14.0% 19.0% 12.0% 18.0%
@@
Qwen2.5-14B-Instruct (+PA) 16.0% 14.0% 24.0% 25.0% 16.0% 6.0% 18.0%
Qwen2.5-14B-Instruct (+NPA) 23.0% 13.0% 24.0% 23.0% 20.0% 10.0% 20.0%
@@
As shown in table 2, both online adaptation (PA) and dynamics grounding (NPA) improves performance across models and tasks. NPA yields substantial improvements for GPT4.1, which is better able to leverage the provided dynamics due to its superior instruction-following abilities. It is also evident in table 2 that our GPT4o-mini surpasses WMA that uses a learned world model by a large margin. For Qwen2.5-14B-Instruct, both methods provide gains, highlighting the general applicability of test-time adaptation. The improvements of non-parametric test-time adaptation is especially significant when using a model with stronger instruction-following ability (i.e., GPT4.1 and GPT4o-mini). We hypothesize this is due to stronger instruction-following abilities of stronger models.
@@
The two proposed strategies are computationally efficient, though they have different profiles. parametric test-time adaptation is highly efficient for real-time use. As shown in table 4, a single update step adds only a 3% latency overhead. In contrast, non-parametric test-time adaptation is a one-time, deployment-time investment designed to be far more lightweight than traditional world-modeling approaches like the WMA baseline. For example, while WMA requires synthesizing 870 tasks to collect trajectories for training a separate Llama-3.1-8B model, our method uses only 50 exploratory rollouts and requires no additional model training. This streamlined process avoids WMA’s costly training phase and its significant inference overhead, where action simulation takes approximately 140 seconds per step. The total upfront cost for our non-parametric procedure on a single WebArena website is approximately 7M tokens, a cost that is amortized across all subsequent tasks. This investment pays significant dividends where environmental dynamics are unpredictable. For instance, on the challenging multi-site split, the success rate of our agent increases from 2% to 23%. This demonstrates the value of efficiently learning an environment’s dynamics at test time to compensate for the limitations of an agent’s pre-training.
@@
The smaller gains for NPA on simpler websites like ”Shopping” suggest that when an environment’s dynamics align with common sense priors (e.g., clicking ’search’ shows a search bar), the explicit dynamics provide less new information, and the overhead of a longer context may slightly hinder performance. In these simple settings, longer input on dynamics may negatively impact the model’s effectiveness, as prior work has shown that excessively long contexts can hurt performance (beltagy_longformer_2020; liu_lost_2024). We further analyze this phenomenon and provide exemplar environment dynamics in section A.5. Furthermore, our test of a naive hybrid approach showed mixed results, underperforming the non-parametric method alone on BFCLv3 (21.0% vs 22.0%). This suggests a simple combination is insufficient and may create interference between the two adaptation signals. A key challenge for future work is developing a more principled integration of semantic guidance and distributional alignment.
Table 5: Ablation on exploration policy and dynamics extractor backbones on WebArena. Here we use different backbones of exploration policy and dynamics extractor. We found for non-parametric adaptation—using the same LLM agent (itself) improves the same as using a stronger one.
@@
GPT4o-mini (baseline) - - 12.0
GPT4o-mini GPT4.1 GPT4.1 18.0 (+6.0)
GPT4o-mini (self improvement) GPT4o-mini GPT4o-mini 19.0 (+7.0)
Qwen2.5-14B-Instruct (baseline) - - 17.0
Qwen2.5-14B-Instruct GPT4.1 GPT4.1 20.0 (+3.0)
Qwen2.5-14B-Instruct (self improvement) Qwen2.5-14B-Instruct Qwen2.5-14B-Instruct 20.0 (+3.0)
@@ 4.3 Ablations
Non-parametric adaptation enables self-improvement
We ablate the choice of backbone LLMs for exploration and dynamics extraction in our non-parametric adaptation method (table 5). Results show that using the same model for both components (self-improvement) performs as well as using a stronger model, indicating that our approach is robust to the choice of exploration policy and dynamics extractor.
@@
We conduct an ablation experiment on BFCLv3, we find that filtering improves success rate across exploration budgets. Specifically, with 10 episodes, success rate rises from 61.0 to 64.0 (+3.0).
Parametric adaptation is robust to different hyperparameters
@@
Larger model benefits from a larger learning rate
We test with a very large learning rate (i.e., learning rate of 1.0), we find that larger models (i.e., 14B and 32B models) improves better with a larger learning rate (i.e., learning rate of 1.0). We find that because the dimension of the adaptation vector for 14B and 32B models is the same (i.e., 5120), and is much larger compared to the 7B model’s (i.e., 3584).

### v4

@@ 4.1 Setup
We evaluate our approach on three realistic agentic benchmarks: BFCLv3 (Patil et al., ), WebArena (Zhou et al., 2024a), and Tau-Bench (Yao et al., 2024), which together cover both web navigation and conversational function calling tasks. As for baselines, we evaluate open-source Qwen2.5-14B-Instruct (Qwen et al., 2025) model and closed-source GPT-4.1. On WebArena, we additionally compare to the World-model-augmented (WMA) agent from (Chae et al., 2025) as a baseline. This approach first collects exploration demonstrations, then trains a Llama-3.1-8B-Instruct (Grattafiori et al., 2024) model to predict the next state given an action (i.e., a learned world model). During evaluation, this world model is used to provide next-state predictions to the agent on-the-fly, helping the agent reason about the consequences of its actions. In this comparison, we use the same GPT-4o mini as the original WMA experiments. Aside from the syntactic alignment and dynamics grounding experiments, we provide ablation studies in section 4.3.
For all dynamics grounding experiments, we leverage GPT-4.1 as an exploration policy to extract environmental dynamics. For each environment, we conduct $10$ exploration episodes to gather a diverse set of dynamics. For environment dynamics filtering, we apply o3222https://platform.openai.com/docs/models/o3 to remove repetitive and trivial environment dynamics. We also include ablation studies on the number of exploration episodes and the effect of filtering in section 4.3.
@@
The WebArena benchmark consists of five self-hosted websites and a total of 812 tasks. Many tasks require interacting across multiple websites, which poses additional challenges for LLM-based agents. For evaluation, WebArena employs a string-matching-based mechanism to assess correctness, providing more stable and objective results compared to benchmarks using LLMs to score, such as (He et al., 2024; Gou et al., 2025). For dynamics grounding, we first synthesize $10$ personas via prompt instruction LABEL:lst:web_persona_synthesis for each website using GPT-4o (OpenAI et al., 2024) and with seed personas from NNetNav (Murty et al., 2025). Then for each persona, we roll-out a GPT-4.1 agent to explore the environment guided by the persona description, with the exploration prompt instruction LABEL:lst:web_exploration derived from NNetNav, and with a maximum exploration budget of 30 steps. To extract environment dynamics from interaction logs, we summarize each state transition pair (observation, action taken, new observation) with prompt instruction LABEL:lst:web_dynamics_extraction, and concatenate all summarized state transition pairs to obtain the list of environmental dynamics $E=\{e\}_{i=1}^{M}$ for each website. Finally, we utilize the o3 model to remove repetitive and trivial environment dynamic entries from the list via prompt instruction LABEL:lst:web_filter_dynamics, resulting in $E_{\text{clean}}$. We provide all prompts we use in appendix A. With syntactic alignment, we reset the adaptation vector after each episode and we use a learning rate of 0.1 and training step of 1 for all experiments. During evaluation, we use the text-based accessibility tree as the observation type and run experiments using BrowserGym [^3] https://github.com/ServiceNow/BrowserGym and AgentLab [^4] https://github.com/ServiceNow/AgentLab with a maximum step of $30$.
Table 2: Main results on WebArena, WebVoyager, BFCLv3, and Tau-Bench benchmarks. We report task success rates (%) for each model and adaptation method. For Tau-Bench, we average runs across 5 seeds and use a custom more stable codebase. Both test-time adaptation strategies improve performance. ”N/A” indicates not applicable—AWM requires training a web-specific world model on state transitions, which is not applicable to Tau-Bench (conversational only) or BFCLv3 (no explicit web-like state transition data); Dynamics grounding does not operate in conversational Tau-Bench as there is no explicit and fixed state transition rules in the environment.
@@
GPT-4.1 30.0 55.5 - -
GPT-4.1 (+ DG) 35.0 (+5.0) 64.0 (+8.5) N/A N/A
GPT-4o mini 12.0 - - -
GPT-4o mini (+ WMA) 13.5 (+1.5) N/A N/A N/A
GPT-4o mini (+ DG) 18.0 (+6.0) - N/A N/A
@@
Qwen2.5-14B-Instruct (+ SA) 18.0 (+1.0) 20.0 (+1.5) 25.2 (+3.6) 44.9 (+1.6)
Qwen2.5-14B-Instruct (+ DG) 20.0 (+3.0) 22.0 (+3.5) N/A N/A
@@
BFCLv3 (Patil et al., ) is a benchmark with various realistic APIs to evaluate the functions calling abilities of LLM. It contains $8$ different API domains. We provide the number of APIs for each environment in table 8. We choose the multi-turn-base split to assess the agent’s multi-turn function calling capabilities. This split is multi-turn and multi-step, meaning that instead of just handling a single tool call in a one-off request, models have to carry out sequences of function calls across multiple dialogue turns. BFCLv3 uses state-based evaluation (checking the system’s resulting state) and response-based evaluation (checking whether the call sequences are valid and minimally sufficient) to assess correctness, therefore being stable and deterministic. For dynamics grounding, we synthesize $10$ exploration tasks with GPT-4o conditioned on the function documentation for each domain via prompt instruction LABEL:lst:bfcl_exploration_synthesis. Concretely, we instruct the LLM to think about what functions it wants to explore in order to find out unpredictable outcomes or interesting behaviors, given the function documentation. We use a GPT-4.1 agent to execute the exploration task and we collect interaction logs for environment dynamics extraction. During exploration, we ask the exploration agent to decide if the steps taken so far are sufficient to answer the exploration task with prompt instruction LABEL:lst:bfcl_exploration_user_prompt. The environment dynamics are extracted using prompt instruction LABEL:lst:bfcl_extract_dynamics. We follow the same procedure as in WebArena to remove repetitive and trivial environment dynamics with prompt instruction LABEL:lst:bfcl_filter_env_dynamics. For syntactic alignment, we reset adaptation vector at each turn (when a new user query arrives).
@@
Tau-Bench (Yao et al., 2024) is a realistic function calling benchmark that simulates real-world service agents interacting with an LLM-simulated user. It comprises two domains: airline (50 tasks) and retail (115 tasks). The airline domain is generally more challenging than retail, as it features a much larger database, increasing the complexity of the tasks. Tau-bench’s use of an LLM-simulated user agent introduces significant variance in evaluation results, as the simulated human may drift from the original instruction or be influenced by the agent’s responses over multiple conversational turns. To address this, we adopt the improved evaluation codebase from (Prabhakar et al., 2025), which substantially reduces reward variance and improves evaluation fidelity. Specifically, we (1) fix incorrect labels, (2) set the temperature of the LLM-simulated user agent to 0 for deterministic responses, and (3) employ a Best-of-N (N=5) sampling strategy combined with a self-critique mechanism, allowing the simulated user to better adhere to the task instruction and avoid being misled by the agent’s outputs. This protocol has been shown to improve consistency in agent performance evaluation across multiple trials according to (Prabhakar et al., 2025). Tau-Bench consists solely of API-based tasks without observable state transitions, meaning there are no environment dynamics to extract. As a result, dynamics grounding, which relies on extracting and leveraging such dynamics, is not applicable to this benchmark. Therefore, we only evaluate the syntactic alignment method on Tau-Bench. To further reduce variance, we report results averaged over 5 random seeds in table 2.
@@ 4.2 Results and Analysis
Table 3: Success rates (%) on the WebArena benchmark across different websites and models. Both adaptation strategies improve performance over baselines. and dynamics grounding (DG) surpasses WMA on GPT-4o mini. DG improves success rate substantially on complex multi-site split.
@@
Model Gitlab Shop Admin Shop Reddit Map Multi Avg
GPT-4.1 48.0% 31.0% 37.0% 29.0% 28.0% 2.0% 30.0%
GPT-4.1 (+DG) 45.0% 27.0% 33.0% 42.0% 31.0% 23.0% 35.0%
GPT-4o mini 9.0% 12.0% 18.0% 8.0% 12.0% 12.0% 12.0%
GPT-4o mini (+WMA) 13.5%
GPT-4o mini (+DG) 18.0% 14.0% 26.0% 14.0% 19.0% 12.0% 18.0%
@@
Qwen2.5-14B-Instruct (+SA) 16.0% 14.0% 24.0% 25.0% 16.0% 6.0% 18.0%
Qwen2.5-14B-Instruct (+DG) 23.0% 13.0% 24.0% 23.0% 20.0% 10.0% 20.0%
@@
As shown in table 2, both syntactic alignment (SA) and dynamics grounding (DG) improve performance across models and tasks. dynamics grounding yields substantial improvements for GPT-4.1, which is better able to leverage the provided dynamics due to its superior instruction-following abilities. It is also evident in table 2 that our GPT-4o mini surpasses WMA that uses a learned world model by a large margin. For Qwen2.5-14B-Instruct, both methods provide gains, highlighting the general applicability of test-time adaptation. The improvements of dynamics grounding are especially significant for models with stronger instruction-following abilities (e.g., GPT-4.1), likely because they can better leverage the provided dynamics context.
@@
The two proposed strategies are computationally efficient, though they have different profiles. syntactic alignment is highly efficient for real-time use. As shown in table 4, a single update step adds only a 3% latency overhead. In contrast, dynamics grounding is a one-time, deployment-time investment designed to be far more lightweight than traditional world-modeling approaches like the WMA baseline. For example, while WMA requires synthesizing 870 tasks to collect trajectories for training a separate Llama-3.1-8B model, our method uses only 50 exploratory rollouts and requires no additional model training. This streamlined process avoids WMA’s costly training phase and its significant inference overhead, where action simulation takes approximately 140 seconds per step. The total upfront cost for our dynamics grounding procedure on a single WebArena website is approximately 7M tokens, a cost that is amortized across all subsequent tasks. This investment pays significant dividends where environmental dynamics are unpredictable. For instance, on the challenging multi-site split, the success rate of our agent increases from 2% to 23%. This demonstrates the value of efficiently learning an environment’s dynamics at test time to compensate for the limitations of an agent’s pre-training.
@@
The smaller gains for DG on simpler websites like “Shopping” suggest that when an environment’s dynamics align with common sense priors (e.g., clicking ’search’ shows a search bar), the explicit dynamics provide less new information, and the overhead of a longer context may slightly hinder performance. In these simple settings, longer input on dynamics may negatively impact the model’s effectiveness, as prior work has shown that excessively long contexts can hurt performance (Beltagy et al., 2020; Liu et al., 2024). We further analyze this phenomenon and provide exemplar environment dynamics in section A.5. Furthermore, our test of a naive hybrid approach showed mixed results, underperforming the dynamics grounding method alone on BFCLv3 (21.0% vs 22.0%). This suggests a simple combination is insufficient and may create interference between the two adaptation signals. A key challenge for future work is developing a more principled integration of semantic guidance and distributional alignment.
Table 5: Ablation on exploration policy and dynamics extractor backbones on WebArena. Here we use different backbones of exploration policy and dynamics extractor. We found for dynamics grounding—using the same LLM agent (itself) improves the same as using a stronger one.
@@
GPT-4o mini (baseline) 12.0
GPT-4o mini GPT-4.1 GPT-4.1 18.0 (+6.0)
GPT-4o mini (self) GPT-4o mini GPT-4o mini 19.0 (+7.0)
Qwen2.5-14B-Instruct (baseline) 17.0
Qwen2.5-14B-Instruct GPT-4.1 GPT-4.1 20.0 (+3.0)
Qwen2.5-14B-Instruct (self) Qwen2.5-14B-Instruct Qwen2.5-14B-Instruct 20.0 (+3.0)
@@ 4.3 Ablations
Dynamics grounding enables self-improvement
We ablate the choice of backbone LLMs for exploration and dynamics extraction in our dynamics grounding method (table 5). Results show that using the same model for both components (self-improvement) performs as well as using a stronger model, indicating that our approach is robust to the choice of exploration policy and dynamics extractor.
@@
We conduct an ablation experiment on BFCLv3 and find that filtering improves success rate across exploration budgets. Specifically, with 10 episodes, success rate rises from 61.0 to 64.0 (+3.0).
Syntactic alignment is robust to different hyperparameters
@@
Larger models benefit from a larger learning rate
When testing with a very large learning rate (i.e., learning rate of 1.0), we find that larger models (i.e., 14B and 32B models) benefit more with a larger learning rate. We attribute this to the fact that the dimension of the adaptation vector for 14B and 32B models is the same (i.e., 5120), and is much larger compared to the 7B model’s (i.e., 3584).